OpenStudy (thatcreepyboy):

Write a proof to show that a rectangle has congruent diagonals. Be sure to create and name the appropriate geometric figures. I need help. Please. I don't have that much time.

jhonyy9 (jhonyy9):

do you can drawing a rectangle ?

OpenStudy (thatcreepyboy):

I don't think so I can try

jhonyy9 (jhonyy9):



OpenStudy (thatcreepyboy):

Wow that's perfect

jhonyy9 (jhonyy9):

ty

OpenStudy (thatcreepyboy):

No problem.

jhonyy9 (jhonyy9):

how you think that you can proving that the diagonales are congruent ?

OpenStudy (thatcreepyboy):

Nope :/ idk how

OpenStudy (lord_box):

What I would do is create two triangles, acknowledge that the two vertical sides are congruent, then use the pythagorean theorem to prove that the two diagonals are congruent

jhonyy9 (jhonyy9):

hope you know that a rectangle has opposite sides paralleles and equal - yes ?

OpenStudy (thatcreepyboy):

Right

jhonyy9 (jhonyy9):

so than you know these and use what above wrote Lord Box so just you need to prove that bd = ac using pythagora"s theorem what say that hypotenuse squared is equal 1 leg squared + 2 leg squared so ac^2 = ad^2 +dc^2 and on the second one triangle will be bd^2 = bc^2 +dc^2 and now using these above wrote that the opposite sides of a rectangle are equal and parallel so from this result that ad=bc and dc = ab so than ac^2 = ad^2 +dc^2 bd^2 = bc^2 +dc^2 and here because ad=bc so we can replace ad in place of bc and will get bd^2 = ad^2 +dc^2 what is exactly the same with the first equation so what mean that these diagonals are congruent hope this helped you and this is understandably easy

OpenStudy (thatcreepyboy):

Thank you I got it!

jhonyy9 (jhonyy9):

welcome was my pleasure good luck bye bye

OpenStudy (thatcreepyboy):

Bye bye!! Have a nice day! :)

jhonyy9 (jhonyy9):

nice night but thank you bc. here is in the night 0,05 oclock - in central Europe - Hungary

jhonyy9 (jhonyy9):

bye

OpenStudy (thatcreepyboy):

..lthats so cool!